OpenStudy (anonymous):

Quadrilateral PQRS is inscribed in circle A. Which statement is necessarily true?

OpenStudy (anonymous):

What are the statements?

OpenStudy (anonymous):

answers m R = m S m R + m S = 180° mR = mS m R + m S = m P + m Q

OpenStudy (anonymous):

@vanessaeasler352

OpenStudy (welshfella):

hint opposite angles of a quadrilateral inscribed in a circle add up to 180 degrees

OpenStudy (anonymous):

I can't see the picture but the opposite angles of PQRS add up to 180

OpenStudy (anonymous):

So if R and S are on opposite corners then adding them together would equal 180

OpenStudy (welshfella):

Yes but r and s are not on opposite sides But you can work the measure of angles r an s

OpenStudy (welshfella):

* work out

OpenStudy (welshfella):

R is opposite p and s is opposite q And p and q both equal 92 degrees

OpenStudy (welshfella):

I would draw it for you but I can't on my iPad

OpenStudy (welshfella):

R = 180 -P s = 180 - q

OpenStudy (welshfella):

So if angles p and q are equal what does that say about r and s?

OpenStudy (welshfella):

I've more or less given you the answer....